# beSnake: A routing algorithm for scalable spin-qubit architectures

### v1

@@ IV The beSnake routing algorithm
The main objective of beSnake is to enable non-adjacent operand qubits of two-qubit gates to interact with each other by strategically shuttling qubits around the topology. It does so by considering multiple shortest paths between the operands and implementing an adjustable path selection heuristic to potentially minimize the added shuttle operations as much as possible. Upon a path selection, beSnake uses a breadth-first search (BFS) exploration for shuttling qubits; hence, the first part of beSnake’s name comes from the first letter of BFS. This approach ensures the minimization of qubit movements, viewed from the qubit operand’s perspective, significantly reducing the pool of qubits considered for movement. In terms of managing parallelized gates, each is routed in turn, while their operands are held in place once they become adjacent until all of the gates are attempted. In case of failed attempts, three strategies are implemented to overcome them, discussed in Section IV.4.1, including a lookback mechanism. In the end, the successful gate attempts can be simultaneously executed, thereby maintaining the circuit depth as low as possible.
In order to further elaborate on the functionalities included in beSnake, we will use three representative cases based on the squared grid topology shown in Figure 0(b). In these examples, we will use the notion of the intermediate representation (IR) for the input circuits as well as output routed circuits with the variable names ”ir_input” and ”ir_output”, respectively. Within the IR, the gates of a cycle are placed inside square brackets, and commas separate each gate. Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates. Moving on, the temporary ”goals” variable contains a single cycle of the input circuit given to beSnake for routing, and the ”goal” variable holds the current gate from ”goals” to be routed. Finally, we will use the ”prev_goal_qubits” variable, which contains the pairs of operands that have already been satisfied from ”goals”.
@@ IV.1 Case 1: Following a free path
The first example illustrates the most simple case of qubit routing. We assume a quantum circuit with only two two-qubit gates (tqg) in the first cycle (see Figure 1(a)), whose interacting qubits are placed in non-neighboring positions (see Fig. 1(b)) and therefore require routing. The corresponding IR representation of the quantum circuit can be derived as ir_input = { [ tqg [1,2], tqg [3,4] ] }. Note that there can be more than one cycle in an IR, but we assume one for illustrative purposes. As depicted in Figure 1(b), the beSnake algorithm takes the first gate of the cycle (variable ”goals”) and assigns it to ”goal”. It then calculates the shortest path between its operands (1 and 2) by using the NetworkX.all_shortest_paths() function. Note that in this case, there is only one shortest path from qubit 1 to qubit 2, comprising just two nodes. More precisely, qubit 1 will need to move upwards two times.
After these movements, qubits 1 and 2 will be adjacent, as shown in Figure 1(c), and they will remain in that position until the two-qubit gate is executed. Note that the ”ir_output” list is updated accordingly, including two new shuttles. In addition, the algorithm updates ”prev_goal_qubits” with qubits 1 and 2, indicating that they will remain in their positions until the rest of the gates are routed. It then proceeds to the next gate in ”goals” and repeats the same process. The outcome, as shown in Figure 1(d), is that the qubits of both gates within ”goals” have been successfully routed, and now the ”ir_output” encompasses all necessary shuttles to position the qubit operands adjacent to each other. After that, the two-qubit gates can be executed together as originally given from the ”ir_input”. At this stage, the router has completed the task, and more optimization passes can further improve the output ”ir_output”, such as a scheduler parallelizing even more the shuttles required for moving qubits 1 and 3.
This example showed a straightforward operation; however, as we will explore in subsequent examples, certain qubits may block the routing path and, therefore, necessitate relocation to allow an operand qubit to progress along the path.
@@ IV.2 Case 2: Following a blocked path
In our second example, we address a similar situation, but this time, qubits are blocking the shortest paths, as shown in Figure 2(a). Utilizing the Networkx.all_shortest_paths() function, we identify four potential shortest paths for addressing the first two-qubit gate between qubits 1 and 2, with each path comprising three steps. For simplicity, we focus only on two paths, labeled as ’a’ and ’b’, as illustrated.
beSnake employs a two-tiered filtering criterion to optimize the path selection. Firstly, it will keep the path(s) with the least obstacle qubits, and out of the remaining paths, it will select one with the highest accumulated degree of traversed nodes. This is to mitigate the likelihood of congestion in the less-connected areas, such as the corners of this topology. This heuristic aims to minimize obstacle encounters and reduce the overhead associated with qubit movement.
Upon comparison, both paths ’a’ and ’b’ involve a single obstacle qubit; however, path ’b’ has notably higher accumulated degree of the traversed nodes, satisfying our selection criteria. When initiating the first step of the preferred shortest path ’b’, as indicated by an orange arrow in Figure 2(b), we encounter a blockage due to qubit 6. Then, an exploration of potential movements for qubit 6 to clear the path is conducted. Three possible directions are identified: moving left towards qubit 5, moving up along the shortest path, and moving to the right. Of these possible movements, only one is practical – moving to the right. By doing so, we avoid the additional future overhead that would arise from moving it upwards.
@@ IV.3 Case 3: Satisfying mutiple two-qubit gates
In the complex scenario depicted in Figure 3(a), beSnake is tasked with satisfying four gates within a cycle amidst a denser qubit topology. By considering routing for these four gates simultaneously, we can achieve a lower depth overhead compared to routing for each separately. We focus on the initial gate listed in ”goals”. As always, we look for the shortest path with the least number of qubit obstacles and with the largest number of adjacent edges. Among the six conceivable paths, each share an identical qubit count; however, only two paths have the most connections to other vertices. The algorithm resolves ties by randomly selecting a path, path ’b’ for this example.
@@
The subsequent phase involves determining the sequence of shuttle movements required to reposition qubit 1 one step further in the chosen shortest path. Figure 3(b) illustrates the exploratory moves beginning from qubit 1. Intuitively, we can envision this as a breadth-first search style of exploration where qubits sequentially push each other until the first vacancy is encountered. Expressly, qubit 1 is set to push qubit 2 (indicated by an orange arrow), while qubit 2 tries to push qubits 4 and 6 (blue arrows), and qubits 4 and 6 try to push qubits 3, 5, and 7 (purple arrows). At this point, a solution has been identified, rendering additional exploration unnecessary, as it would certainly increase the number of moves.
@@
Figure 3(c) updates ”ir_output” with the first sequence of moves from the first step of qubit 1, within a single cycle. This figure also displays the exploration of qubit 1’s next move, with each expansion layer marked by arrows colored orange, blue, and purple, respectively. At this point, we see two possible solutions, moving qubits 3 and 5. Both solutions need the same number of shuttles. beSnake then becomes noise-aware by selecting the one with the highest accumulated shuttle fidelity based on predefined fidelity attributes for each coupling connection. The accumulated fidelity is calculated by multiplying the fidelities of each shuttle on the specific locations. This mechanism is used in case different fidelities are associated with each coupling link, thus maximizing the circuit’s fidelity. As a result, the movement of qubit 3 is chosen for this example, and in Figure 3(d), the sequence of shuttles is displayed in the second cycle of ”ir_output”. In the same figure, we have fast forward to the final step of the shortest path and see the three cycles within ”ir_output” and updates to ”prev_goal_qubits”. The ”goal” now includes the next gate from ”goals” and the previous gate has been removed. A single shortest path is discovered since it is not allowed to push fixed qubits, yet qubit 6 is immobilized. This is because the operand qubits are prohibited from being pushed by any other qubit as well. Consequently, beSnake omits this gate and proceeds to the subsequent one. This mechanism is one of three mechanisms implemented (see section IV.4.1) to deal with unsuccessful attempts when routing certain two-qubit gates. In this particular case, we will move on to route the next available gate in ”goals”, the tqg [0,7], which might potentially move qubits around the topology and free the blockage shown in Figure 3(d).
@@
Figure 3(e) shows that the next goal is immediately satisfied, and the algorithm reattempts the routing of the previously unsuccessful gate, the tqg [5,2]. This is the lookback mechanism used to iterate over the remaining gates in ”goals”. Once tqg [7,0] is accomplished, it is removed from ”goals”, and the algorithm revisits the ”goals” set, starting with any gates that previously failed. However, the tqg [5,2] remains unsuccessful due to the same reasons. Figure 3(f) demonstrates the resolution of the following gate, tqg [6,3], the next gate in ”goals”, at which stage all gates have been satisfied except for one. A repeated attempt yields no solution, as no shortest path is viable anymore between qubits 5 and 2. Due to this fact, we update ”ir_output” by scheduling only the satisfied gates, thus splitting the original cycle.
@@
Subsequently, all qubits in ”prev_goal_qubits” are released, and the algorithm reattempts the unsolved gate. This time, a solution is possible, and Figure 4(a) depicts the successful outcome.
@@ IV.4 Special cases
This section introduces the mechanisms employed by beSnake to address special routing cases. We outline three primary mechanisms beSnake utilizes to circumvent blockades as in Figure 3(c), showcasing the algorithm’s potential to adapt in conflicts.
@@ IV.4.1 Mechanisms for dealing with blockades
An obstacle qubit might disallow the movement of any operand qubit within the shortest path. For example, in Figure 3(c), obstacle qubit 6 is blocking the very first step in the shortest path. However, there can be cases where a similar situation is reached at any step in the path due to qubits being pushed in areas of the topology with fewer connections (edges), such as the corners of squared topologies. Then, the following mechanisms will be tried in this order:
@@
Revisit the same gate (”goal”) after satisfying other ones. This lookback mechanism was used in Figure 3(e).
@@
Since the first two mechanisms are demonstrated in Section IV.3, we are now focusing on the third; If we isolate the scenario in Figure 3(d) and suppose the first two mechanisms failed, then we will insert a forced SWAP between qubit 5 and 2 to overcome this obstacle. This mechanism was implemented to efficiently resolve such rare scenarios where qubit obstacle(s) are located in between both operands while having no available edges to be pushed away. This can be triggered in topologies that may have only two coupling links at one or more quantum dots. In the case of a square grid, this place is at the corners. Alternatively, this can be addressed by taking another path, not necessarily a shortest path, that avoids such restricted regions. However, this does not guarantee a solution and most likely will result in more overhead. In Section VI, more alternatives are discussed.
@@ IV.4.2 Handling shuttle-based Z rotations
Shuttle-based Z rotation gates are a unique high-fidelity implementation of a Z rotation with two time-sensitive qubit shuttles to and from a neighboring column in either direction [1, 49, 50]. beSnake can determine which direction (left or right) will result in the least sequence of shuttles. Then, it will fix the positions (one empty and one containing the qubit operand of the Z gate) so that the return shuttle can be done immediately. To demonstrate that, we suppose a shuttle-based Z rotation on qubit 4 in the example of Figure 3(a). This will result in the example shown in Figure 5(a), where moving to the right takes fewer moves than moving left. In Figure 5(b), the original (now empty) location is not allowed to change until the return shuttle is executed. More specifically, if the empty location gets occupied before the return shuttle, qubit 4 cannot return in time. In case there are parallel Z gates to be satisfied in a cycle, beSnake will follow the same logic as described in Figures 3(e) and 3(f), and iteratively try all of them. An exception is applied here where succeeding Z gates can override fixed empty locations from preceding Z gates. Since implementing Z gates involves a second shuttle back to the original location, multiple of these shuttles can take place simultaneously to complete the cycle even if they previously occupied fixed empty locations of other Z gates. One example of this exception is when a shuttle-based Z gate for qubit 3 is scheduled next in the example of Figure 5(b) during the same cycle. Qubit 3 will shuttle to the right since this is the only available direction, overriding the fixed position created by the Z gate of qubit 4. Then, supposing the cycle finishes here, both qubits can return to their original locations with the return shuttle, completing their Z rotations. This mechanism provides better parallelization as more room is accounted for more Z gates while maintaining the same gate overhead (as we would have without this exception). Finally, in case there are Z and two-qubit gates in a cycle, the two-qubit gates will be sorted at the beginning of ”goals” to be tried first. This is because routing for two-qubit gates will more likely require multiple steps to complete, and therefore, in the beginning, it is more important to have as few fixed positions as possible (i.e., more free space to push/move qubits). In Section VI, more ideas are discussed related to sorting the cycle.
@@ IV.5 Optional functionalities
SWAP replacement: beSnake can become more noise-aware by replacing a sequence of shuttles facilitating a step within the shortest path by a SWAP operation, for example, those inserted in Figure 3(c). It will do so when the accumulated fidelity of those shuttles exceeds that of the SWAP, particularly in that location.
Shortest-path optimizations: As described in the previous sections, beSnake will assess multiple shortest paths and heuristically pick one that will most likely produce the least amount of gate and depth overhead. However, this ability is by far the most time-consuming task of beSnake as the build-in all_shortest_paths() function of Networkx, and subsequently, the ”filtration” of all paths is iterated several times. To mitigate that, there is an option to place a time limit on the discovery of the shortest paths, thus substantially reducing the time this process takes. Additionally, there is an option to take a single shortest path by utilizing the Networkx.shortest_path() function. Later in Section V.2.1, we will test these options and determine the ratio between the increased overhead and reduced routing time.

### v2

@@ IV The beSnake routing algorithm
The main objective of beSnake is to enable non-adjacent operand qubits of two-qubit gates interact with each other by strategically shuttling qubits around the topology. It does so by considering multiple shortest paths between the operands and subsequently implementing an adjustable path selection heuristic to potentially minimize the added movements as much as possible. Upon a path selection, beSnake uses a breadth-first search (BFS) exploration for shuttling qubits; hence, the first part of beSnake’s name comes from the first letter of BFS. This approach with the particular shortest path selection ensures the minimization of shuttle operations, viewed from the operand’s perspective, significantly reducing the pool of qubits considered for movement. In terms of managing parallelized gates, each is routed in turn, while their operands are held in place once they become adjacent until all of the gates are attempted. In case of failed attempts, three strategies are implemented to overcome them, discussed in Section IV.4.1, including a lookback mechanism. In the end, the gates can be simultaneously executed, thereby maintaining the circuit depth as low as possible.
In order to further elaborate on the functionalities included in beSnake, we will use three representative cases based on the squared grid topology shown in Fig. 1(b). In these examples, we will use the notion of the intermediate representation (IR) for the input circuits as well as output routed circuits with the variable names ir_input and ir_output, respectively. Within the IR, the gates of a cycle are placed inside square brackets, and commas separate each gate. Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates. Moving on, the temporary goals variable contains a single cycle of the input circuit given to beSnake for routing, and the goal variable holds the current gate from goals to be routed. Finally, we will use the prev_goal_qubits variable, which contains the pairs of operand qubits that have already been satisfied from goals.
@@ IV.1 Case 1: Following a free path
The first example illustrates the most simple case of qubit routing. We assume a quantum circuit with only two two-qubit gates (tqg) in the first cycle (see Fig. 2(a)), whose interacting qubits are placed in non-neighboring positions (see Fig. 2(b)) and therefore require routing. The corresponding IR representation of the quantum circuit can be derived as ir_input = { [ tqg [1,2], tqg [3,4] ] }. Note that there can be more than one cycle in an IR, but we assume one for illustrative purposes. As depicted in Fig. 2(b), the beSnake algorithm takes the first gate of the cycle (variable goals) and assigns it to goal. It then calculates all possible shortest paths between its operands (1 and 2). In this case, there is only one shortest path from qubit 1 to qubit 2, comprising just two nodes. More precisely, qubit 1 will need to move upwards two times.
After these movements, qubits 1 and 2 will be adjacent, as shown in Fig. 2(c), and they will remain in that position until the two-qubit gate is executed. Note that the ir_output list is updated accordingly, including two new shuttles. In addition, the algorithm updates prev_goal_qubits with qubits 1 and 2, indicating that they will remain fixed in their current positions until the rest of the gates are routed. It then proceeds to the next gate in goals and repeats the same process. The outcome, as shown in Fig. 2(d), is that the qubits of both gates within goals have been successfully routed, and now the ir_output encompasses all necessary shuttles to position the qubit operands adjacent to each other. After that, the two-qubit gates can be executed together as originally given from the ir_input. At this stage, beSnake has completed the task, and more optimization passes can further improve the output ir_output, such as a scheduler parallelizing even more the shuttles required for moving qubits 1 and 3.
This example showed a straightforward operation; however, as we will explore in subsequent examples, certain qubits may block the routing path and, therefore, necessitate relocation to open the way.
@@ IV.2 Case 2: Following a blocked path
In our second example, we address a similar situation, but this time, qubits are blocking the shortest paths, as shown in Fig. 3(a). We can identify four potential shortest paths for addressing the first two-qubit gate between qubits 1 and 2, with each path comprising three steps. For simplicity, we focus only on two paths, labeled a and b, as illustrated.
beSnake employs a two-tiered filtering criterion to optimize the path selection. Firstly, it will keep the path(s) with the least obstacle qubits, and out of the remaining paths, it will select one with the highest accumulated degree of traversed nodes. This is to mitigate the likelihood of congestion in the less connected areas, such as the corners of the current topology. This heuristic aims to minimize obstacle encounters and to reduce the overhead associated with qubit movement.
Upon comparison, both paths a and b involve a single obstacle qubit; however, path b has notably higher accumulated degree of the traversed nodes, satisfying our selection criteria. When initiating the first step of the preferred shortest path b, as indicated by an orange arrow in Fig. 3(b), we encounter a blockage due to qubit 6. Then, an exploration of potential movements for qubit 6 to clear the path is conducted and three possible directions are identified: moving left towards qubit 5, moving up along the shortest path, and moving to the right. Of these possible movements, only one is practical – moving to the right. By doing so, we avoid the additional overhead that would otherwise arise later by moving it again away from the shortest path.
@@ IV.3 Case 3: Satisfying mutiple two-qubit gates
In the complex scenario depicted in Fig. 4(a), beSnake is tasked with satisfying four gates within a cycle amidst a denser qubit topology. By considering routing for these four gates simultaneously, we can achieve a lower depth overhead compared to routing for each separately. We focus on the initial gate listed in goals. As always, we look for the shortest path with the least number of qubit obstacles and with the largest number of adjacent edges. Among the six conceivable paths, each shares an identical qubit count; however, only two paths have the most connections to other vertices. The algorithm resolves ties by randomly selecting a path, path b for this example.
@@
The subsequent phase involves determining the sequence of shuttle movements required to reposition qubit 1 one step further in the chosen shortest path. Fig. 4(b) illustrates the exploratory moves beginning from qubit 1. Intuitively, we can envision this as a BFS style of exploration where qubits sequentially push each other until the first vacancy is encountered. Expressly, qubit 1 is set to push qubit 2 (indicated by an orange arrow), while qubit 2 tries to push qubits 4 and 6 (blue arrows), and qubits 4 and 6 try to push qubits 3, 5, and 7 (purple arrows). At this point, a solution has been identified, rendering additional exploration unnecessary, as it would certainly increase the number of moves.
@@
Fig. 4(c) updates ir_output with the first sequence of moves from the first step of qubit 1, within a single cycle. This figure also displays the exploration of qubit 1’s next move, with each expansion layer marked by arrows colored orange, blue, and purple, respectively. At this point, we see two possible solutions, moving qubits 3 and 5. Both solutions need the same number of shuttles. beSnake then becomes noise-aware by selecting the one with the highest accumulated shuttle fidelity based on predefined fidelity attributes for each coupling connection. The accumulated fidelity is calculated by multiplying the fidelities of each shuttle on the specific locations. This mechanism is used in case different fidelities are associated with each coupling link, thus maximizing the routed circuit’s fidelity. As a result, the movement of qubit 3 is chosen for this example, and in Fig. 4(d), the sequence of shuttles is displayed in the second cycle of ir_output. In the same figure, we have fast forward to the final step of the shortest path and see the three cycles within ir_output and updates to prev_goal_qubits. The goal now includes the next gate from goals and the previous gate has been removed. A single shortest path is discovered since it is not allowed to push fixed qubits, yet qubit 6 is immobilized. This is because the operand qubits are prohibited from being pushed by any other qubit as well. Consequently, beSnake omits this gate and proceeds to the subsequent one. This mechanism is one of three mechanisms implemented (see section IV.4.1) to deal with unsuccessful attempts when routing certain two-qubit gates. In this particular case, we will move on to route the next available gate in goals, the tqg [0,7], which might potentially move qubits around the topology and free the blockage shown in Fig. 4(d).
@@
Fig. 4(e) shows that the next goal is immediately satisfied, and the algorithm reattempts the routing of the previously unsuccessful gate, the tqg [5,2]. This is the lookback mechanism used to iterate over the remaining gates in goals. Once tqg [7,0] is accomplished, it is removed from goals, and the algorithm revisits the goals set, starting with any gates that previously failed. However, the tqg [5,2] remains unsuccessful due to the same reasons. Fig. 4(f) demonstrates the resolution of the following gate, tqg [6,3], the next gate in goals, at which stage all gates have been satisfied except for one. A repeated attempt yields no solution, as no shortest path is viable anymore between qubits 5 and 2. Due to this fact, we update ir_output by scheduling only the satisfied gates, thus splitting the original cycle.
@@
Subsequently, all qubits in prev_goal_qubits are released, and the algorithm reattempts the unsolved gate. This time, a solution is possible, and Fig. 5(a) depicts the successful outcome.
@@ IV.4 Special cases
This section introduces the mechanisms employed by beSnake to address special routing cases. We outline three primary mechanisms utilized to circumvent blockades showcasing the algorithm’s potential to adapt to conflicts.
@@ IV.4.1 Mechanisms for dealing with blockades
An obstacle qubit might disallow the movement of any operand qubit within the shortest path. For example, in Fig. 4(c), obstacle qubit 6 is blocking the very first step in the shortest path. However, there can be cases where a similar situation is reached at any step in the path due to qubits being pushed in areas of the topology with fewer connections, such as the corners of squared topologies. Then, the following mechanisms will be tried in this order:
@@
Revisit the same gate (goal) after satisfying other ones. This lookback mechanism was used in Fig. 4(e).
@@
Since the first two mechanisms have been demonstrated in Section IV.3, we are now focusing on the third one; If we isolate the scenario in Fig. 4(d) and suppose the first two mechanisms failed, then we will insert a forced SWAP between qubit 5 and 2 to overcome this obstacle. This mechanism was implemented to efficiently resolve such rare scenarios where qubit obstacle(s) are located in between both operands while having no available edges to be pushed away. This can be triggered in topologies that may have only two coupling links at one or more nodes (quantum dots). In the case of a square grid, this place is at the corners. Alternatively, this can be addressed by taking another path, not necessarily a shortest path, that avoids such restricted regions. However, this does not guarantee a solution and most likely will result in more overhead. In Section VI, more alternatives are discussed.
@@ IV.4.2 Handling shuttle-based Z rotations
Shuttle-based Z rotation gates are a unique high-fidelity implementation of a Z rotation with two time-sensitive qubit shuttles to and from a neighboring column in either direction [1, 52, 53, 25]. This kind of single-qubit manipulation technique, otherwise called “hopping spins” [54], can, in fact, achieve an arbitrary rotation axis, but in this work, we assume the proposal in [25] for simplicity. beSnake can determine which direction (left or right) will result in the least sequence of shuttles. Then, it will fix the positions (one empty and one containing the qubit operand of the Z gate) so that the return shuttle can be inserted immediately. To demonstrate that, we suppose a shuttle-based Z rotation on qubit 4 in the example of Fig. 4(a). This will result in the example shown in Fig. 6(a), where moving to the right takes fewer moves than moving left. In Fig. 6(b), the original (now empty) location of qubit 4 is not allowed to change until the return shuttle is executed. More specifically, if the empty location gets occupied before the return shuttle, qubit 4 cannot return in time. In case there are parallel Z gates to be satisfied in a cycle, beSnake will follow the same logic as described in Figures 4(e) and 4(f), and iteratively try all of them. An exception is applied here where succeeding Z gates can override fixed empty locations from preceding Z gates. Since implementing Z gates involves a second shuttle back to the original location, multiple of these shuttles can take place simultaneously to complete the cycle even if they previously occupied fixed empty locations of other Z gates. One example of this exception is when a shuttle-based Z gate for qubit 3 is scheduled together with the Z gate on qubit 4 in Fig. 6(b). Qubit 3 will shuttle to the right since this is the only available direction, overriding the fixed position created by the Z gate of qubit 4. Then, supposing the cycle finishes here, both qubits can return to their original locations with the well-timed return shuttles, completing their Z rotations. This exception provides better parallelization as extra room is provided for more Z gates while maintaining the same gate overhead (as we would have without this exception).
Finally, in case there are Z and two-qubit gates in a cycle, the two-qubit gates will be sorted at the beginning of goals to be tried first. This is because routing for two-qubit gates will more likely require multiple steps to complete, and therefore, in the beginning, it is more important to have as few fixed positions as possible (i.e., more free space to push/move qubits). In Section VI, more ideas are discussed related to sorting gates in cycles.
@@ IV.5 Optional functionalities
SWAP replacement: beSnake can become even more noise-aware by replacing a sequence of shuttles facilitating a step within the shortest path by a SWAP operation. One example is the sequence of shuttles inserted in Fig. 4(c). It will do so when the accumulated fidelity of those shuttles exceeds that of the SWAP, particularly in these locations.
Shortest-path optimizations: As described before, beSnake will assess multiple shortest paths and heuristically pick one that will most likely produce the least amount of gate and depth overhead. However, this ability is by far the most time-consuming task of beSnake as the discovery of all shortest paths, and subsequently, their filtration is iterated several times. To mitigate that, there is an option to place a time limit on the discovery of the shortest paths, thus substantially reducing the time this process takes. Additionally, there is an option to take a single shortest path, which significantly speeds up the process. Later in Section V.2.1, we will deliberately test these options and determine the routing time benefits at the expense of increased circuit overhead.